Fixed costs remain constant at $400,000 per month. During high-output months variable costs are $320,000, and...

  1. Fixed costs remain constant at $400,000 per month. During high-output months variable costs are $320,000, and during low-output months variable costs are $80,000. What are the respective high and low indirect-cost rates if budgeted professional labor-hours are 16,000 for high-output months and 4,000 for low-output months?
  1. $56.20 per hour; $120.00 per hour
  2. $45.00 per hour; $120.00 per hour
  3. $25.00 per hour; $20.00 per hour
  4. $45.00 per hour; $45.00 per hour
  5. None of the above
Answer #1

Total cost at high volume = 400000+320000 = $720000

Total cost at low volume = 400000 + 80000 = $480000

Cost per hour = 720000/16000 = $45 per hour

= 480000/4000 = $120 per hour

Option b. is correct answer.